# Adaptive Depth Networks with Skippable Sub-Paths

### v1

@@ 4.7 Object Detection and Instance Segmentation
5 Conclusions
We propose an architectural pattern for adaptive depth networks and its training method that renders selected sub-paths of the network to incur minimal performance degradation even if they are skipped during inference. Our approach does not train a fixed set of sub-networks iteratively, and, hence significantly shorter training time is required than previous adaptive networks. However, once deployed on devices, it can instantly construct sub-networks of various depths to provide various accuracy-efficiency trade-offs in a single network. We show formally that the proposed architectural pattern and training method reduces overall prediction errors while minimizing the impact of skipping sub-paths. We also empirically demonstrate the generality and effectiveness of our approach using representative convolutional neural networks and vision transformers.

### v2

@@ B.1 Object Detection and Instance Segmentation
B.2 Visual Analysis of Sub-Paths
To investigate how our training method affects feature representations in the mandatory and the skippable sub-paths, we visualize the activation of 3rd residual stage of ResNet50-ADN using Grad-CAM [63]. The 3rd residual stage of ResNet50-ADN has 6 residual blocks and the last three blocks are skippable. In Figure 8-(a), the activation regions of original ResNet50 changes gradually across all consecutive blocks. In contrast, in Figure 8-(b), ResNet50-ADN(FFFF), or super-net, manifests very different activation regions in two sub-paths. In the first three residual blocks, we can observe lots of hot activation regions in wide areas, suggesting active learning of new level features. In contrast, significantly less activation regions are found in the skippable last three blocks and they are gradually concentrated around the target object, demonstrating the refinement of learned features. While ResNet50-ADN(TTTT), or base-net, shares the parameters with the super-net in the first 3 mandatory blocks, their activation regions are very different from the super-net’s. This is because while the super-net and the base-net share parameters in the non-skippable mandatory blocks, they use different batch normalization operators in the mandatory sub-paths. Further, in Figure 8-(c), we can observe that the final activation map of the base-net is very similar to the super-net’s final activation map in Figure 8-(b). This implies that they have similar distributions for the same inputs, as suggested in Section 3.3.
(a) Original ResNet50
(b) ResNet50-ADN (FFFF), or super-net
(c) ResNet50-ADN (TTTT), or base-net
Figure 8: Class Activation Maps of the 3rd residual stages of ResNet50s. (a) Original ResNet50’s activation regions change gradually across all blocks. (b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features. In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target. (c) Even though parameters are shared, the activation map of base-net is very different from super-net’s since they use different batch normalization operators.